OpenStudy (anonymous):

write an equation of the line containing the given point and parallel to the given line. express your answer in the form y=mx+b (-2,6);6x=5y+8

jimthompson5910 (jim_thompson5910):

6x=5y+8 6x-5y=8 ----------- 6x-5y=C 6(-2)-5(6)=C -12-30=C -42=C C=-42 So the parallel equation is 6x-5y=-42. Now solve for y 6x-5y=-42 -5y=-6x-42 y=6/5x+42/5 Answer: \[y=\frac{6}{5}x+\frac{42}{5}\]

OpenStudy (anonymous):

how about (2,6);4x+y=7

jimthompson5910 (jim_thompson5910):

4x+y=C 4(2)+6=C 14=C Parallel equation is 4x+y=14 Solve for y 4x+y=14 y=-4x+14 Answer: y = -4x+14

OpenStudy (anonymous):

oh, it must be perpendicular to the given line

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

do you know the perpendicular equation?

jimthompson5910 (jim_thompson5910):

yes one sec

OpenStudy (anonymous):

thank you

jimthompson5910 (jim_thompson5910):

Answer is \[y=\frac{1}{4}x+\frac{11}{2}\]

OpenStudy (anonymous):

same thing, perpendicular (9,-3);9x+2y=7

jimthompson5910 (jim_thompson5910):

Answer is \[y=\frac{2}{9}x-5\]

OpenStudy (anonymous):

In 1992, the life expectancy of males in a certain country was 69.8 years. In 1996, it was 72.4 years. Let E represent the life expectancy in the year t and let t represent the number of years since 1992. The linear function E(t) that fits the data is.. E(t)=?t+?

jimthompson5910 (jim_thompson5910):

E(t) = 0.65 t+69.8

OpenStudy (anonymous):

round to the nearest tenth

jimthompson5910 (jim_thompson5910):

E(t) = 0.7 t+69.8

jimthompson5910 (jim_thompson5910):

that's of course after rounding to the nearest tenth

OpenStudy (anonymous):

use the function to predict the life expectancyy of males in 2008

OpenStudy (anonymous):

E(16)=

jimthompson5910 (jim_thompson5910):

E(16)=81

OpenStudy (anonymous):

the table list data regarding the average salaries of professional athletes in the years 1991 and 2001 1991 262,000 2001 1,390,000 a: use the data points to find a linear function that fits the data b:use the function to predict the average salary in 2005 and 2010 S(x)=?

jimthompson5910 (jim_thompson5910):

S(x)=112800x+262000

OpenStudy (anonymous):

incorrect

jimthompson5910 (jim_thompson5910):

does it want x to start at 1991?

jimthompson5910 (jim_thompson5910):

oh and is S(x) in terms of thousands?

OpenStudy (anonymous):

Its cool man, I thank you for all the help! The test just ended ofically at 12am mst

jimthompson5910 (jim_thompson5910):

ok my guess is that x was supposed to start at 1991 instead of 0. oh well

OpenStudy (anonymous):

lol thanks